OpenStudy (anonymous):

The sum of 7 and w is greater than or equal to 28 .

OpenStudy (anonymous):

7 + w is greater than or equal to 28

OpenStudy (anonymous):

7 + w > 28 w > 28 - 7 w > 21

OpenStudy (anonymous):

\[7+w \ge 28,w \ge 28-7 or w \ge 21\]